OpenStudy (anonymous):

A village fair offers three throws for $1 and seven throws for $2. If with 187 throws $57 is made, how many people bought three throws

OpenStudy (anonymous):

we need a variable for either the $1 throws or the $2 throws so we can solve this for one of them which one do you pick and what variable?

OpenStudy (anonymous):

could it be x and y

OpenStudy (anonymous):

you want to use two variables that is find also unnecessary but fine

OpenStudy (anonymous):

ok that was wrong how can 187 throws bring in only $57?

OpenStudy (anonymous):

oooh i got that totally wrong doh sorry

OpenStudy (anonymous):

if you only use one variable how would that turn out

OpenStudy (anonymous):

lets see if we can do it this way first

OpenStudy (anonymous):

the total is 187 so \[x+y=187\]

OpenStudy (anonymous):

and also \[\frac{x}{3}+\frac{2y}{7}=57\]

OpenStudy (anonymous):

you can solve that for \(x\) and \(y\) probably easier to multiply the second equation by 21 and start with \[x+y=87\\ 7x+6y=1179\]

OpenStudy (anonymous):

i get \(x=657\) but that is the number of throw, not the number of people the number of people would be one third of that, since they got three throws each

OpenStudy (anonymous):

it says the number of people is suppose to be 25

OpenStudy (anonymous):

really? guess i screwed that one up

OpenStudy (anonymous):

let me see if i can unscrew it if x is the number of people with 3 throws and y with 7 then maybe we should have \[3x+7y=187\]

OpenStudy (anonymous):

and also \[x+2y=57\]

OpenStudy (anonymous):

seems to work nicely, get \(x=25,y=16\)